# Anyon braiding and telegraph noise in a graphene interferometer

### v1

@@ Abstract
The search for anyons, quasiparticles with fractional charge and exotic exchange statistics, has inspired decades of condensed matter research. Quantum Hall interferometers enable direct observation of the anyon braiding phase via discrete interference phase jumps when the quasiparticle number changes. Here, we observe the universal anyonic braiding phase in both the $\nu=1/3$ and $4/3$ fractional quantum Hall states by probing three-state random telegraph noise (RTN) in real-time. We find that the observed RTN stems from anyon quasiparticle number $n$ fluctuations and reconstruct three Aharonov-Bohm oscillation signals phase shifted by $32\pi/$ 3, corresponding to the three possible interference branches from braiding around $n$ (mod 3) anyons. Hence, we fully characterize the anyon exchange statistics using new methods that can readily extend to non-abelian states.
@@ Introduction
FQH Fabry-Pérot (FP) interferometers enable direct measurements of the anyon braiding phase of quasiparticles around a confined cavity [11, 12]. By partially backscattering current at two quantum point contact (QPC) constrictions, the conductance through the FP cavity modulates with the phase accumulated by edge-traveling quantum Hall (QH) quasiparticles [5, 13]. When the number of localized cavity quasiparticles encircled by the interfering edge changes, the resulting phase shift in the interference signal reveals the anyon braiding phase [14]. However, since the observed discrete phase jumps represent charging events in the cavity, resultant Coulomb effects on the interference path can pose complications in probing the braiding effect. In particular, the Coulomb coupling between the bulk and edge must be accounted for when extracting the anyon braiding phase from a small number of phase jump events [15]. Consequently, mitigating Coulomb coupling while maintaining electrostatic tunability has been a major challenge in semiconductor-based interferometers [16, 14].
@@
Figure 1: Graphene FQH interferometer telegraph noise. (A) False-color scanning electron microscope image of a representative device depicting an interference measurement with a single edge channel. We measure the conductance $G=I_{d}/(V_{D}^{+}-V_{D}^{-})$ where $I_{d}$ is the measured drain current and $V_{D}^{\pm}$ are the measured voltages at the indicated locations. The top graphite gate is etched into 8 separately tunable regions controlled by voltages $V_{LG}$, $V_{LSG}$, $V_{PG}$, $V_{MG}$, $V_{RSG}$, and $V_{RG}$. Additionally, voltages $V_{LBG/RBG}$ applied to suspended bridges above the two QPCs independently set their transmissions. See Fig. S1 for more details. (B) Hall conductivity $\sigma_{xy}$ and longitudinal conductivity $\sigma_{xx}$ measured with contacts in the left reservoir of the device highlighting the robust plateaus at $\nu=0$, $1/3$, $1$, and $4/3$, which are the quantum Hall states we use in this work. (C) Derivative of $\sigma_{xy}$ showing the expected magnetic field dependence of the well-developed quantum Hall states. (D) Conductance $G$ across the left/right QPC tuned by $V_{LBG/RBG}$ with the right/left QPC deactivated and the bulk in $\nu=4/3$. See Fig. S2 for more details on QPC tunings. Inset: schematics of the edge channel configuration and tunneling at a QPC. As $V_{LBG/RBG}$ increases, the QPC opens more and transmits the outer integer edge (pink) first and eventually transmits the inner fractional edge (blue) with weak backscattering. (E) Three-state RTN in the conductance $G$ when the fractional edge of $\nu=4/3$ is weakly backscattering (corresponding to the tunings marked by the black/purple dots for the left/right QPC in D). Top inset: cartoons of the three possible interference branches from braiding around $n$ (mod 3) anyons, corresponding to the three conductance levels. Right inset: histogram of the conductance over time with 1000 bins. (F) Extracted sinusoidal oscillations from histograms over approximately 8 minutes for each $V_{PG}$ voltage. Each data point shows the central conductance value of a gaussian fit to a histogram peak, shaded by the total counts summed under the gaussian. See Fig. S3 for similar data in $\nu=1/3$. All data is at $T=20$ mK and $B=12$ T unless otherwise noted.
@@
Graphene-based interferometers offer several advantages for realizing anyon braiding, both for the abelian non-abelian cases. First, both QPCs [17, 18] and interferometers [19, 20, 21, 22, 23] exhibit robust electrostatic tunability for integer QH states. Second, graphite gates reduce Coulomb coupling to reveal Aharonov-Bohm (AB) oscillations [20], which previously required complex quantum well structures in GaAs [16, 14]. Simultaneously, these atomically-flat gates enhance FQH states by screening out charge disorder [24, 25] while allowing precise control of QPC transmission [20, 26, 27] and FP cavity filling factor [28]. Encouragingly, bilayer graphene has displayed a multitude of even-denominator FQH states which may host non-abelian anyons with relatively large energy gaps [29, 30, 31, 32, 33], as required for topologically-protected qubits. Less explored even-denominator states have also been observed in monolayer graphene [34, 35].
@@
Measuring how the magnetic field $B$ changes the total interference phase is crucial in determining if the interferometer falls into the AB or Coulomb-dominated regime [12, 5]. We start for comparison with the $\nu=1$ integer QH state. By repeatedly sweeping the plunger gate, we observe the expected single-sinusoidal oscillations devoid of phase jumps (Fig. 2A). As we change $B$, we observe magnetic-field dependent AB interference (Fig. 2B), where the magnetic field period $\Delta B_{1}$ yields an area $\phi_{0}/\Delta B_{1}=0.80\ \mu\text{m}^{2}$, as expected for interference of electrons in $\nu=1$. A similar measurement in $\nu=1/3$ with varying $B$ and $V_{PG}$ produces an irregular 2D map (Fig. 2C) since repeated plunger gate sweeps are not consistent due to the frequent RTN fluctuations. However, we observe that each plunger gate scan contains stochastic switching between 3 sinusoidal branches, again shifted by $32\pi/$ 3 (Fig. 2D). By superimposing 100 of these sweeps and plotting the histogram of total recorded conductance values over $V_{PG}$, we are able to sample enough instances of the AB oscillation signal of each branch to recreate the three interwoven sinusoids seen in the previous measurement technique discussed in Fig. 1. Fig. 2E exhibits the three interwoven branches at different magnetic fields. As the magnetic field decreases (from the top to bottom panels), we observe that each branch continuously drifts in the increasing $V_{PG}$ direction, creating a negative slope of constant phase (dashed line) analogous to the integer case of Fig. 2B. For $\nu=1/3$, the corresponding flux super-periodicity $\Delta B_{1/3}$ (the magnetic field period for a single branch to return to its initial $V_{PG}$ coordinate) yields $3\phi_{0}/\Delta B_{1/3}=0.83\ \mu\text{m}^{2}$. This value is what one expects for interfering $e/3$ quasiparticles such that $\phi_{0}\rightarrow h/(e/3)=3\phi_{0}$. We similarly observe AB interference and a flux super-periodicity on the fractional edge of $\nu=4/3$ (Fig. S4 in SM), while the integer edge shows typical electron AB interference with no observable RTN.
@@
The striking behavior of the anyon branch switching rate and amplitude along the plateau reveals that the quasiparticle fluctuations and RTN is closely correlated with the compressibility of the FQH state. A natural explanation would be to account for the increase in switching rate with an increase in the number of localized states, and thus anyon trapping sites, near the edges of the FQH plateau. However, a simple increase in the individual switching rates of a constant number of traps across the plateau cannot be ruled out with just this observation alone.
@@
In addition to the decrease in visibility, the RTN also shows an increase in switching rate as temperature increases. Fig. 4C shows temperature dependent $\tau_{s}^{-1}$ for two representative data sets, one at the center (red) and the other at the edge (yellow) of the plateau. While switching rate indeed increases quickly at the high temperature limit, there is a dramatic flat saturation of $\tau_{s}^{-1}$ for both data sets at low temperatures, below about $100$ mK and $50$ mK for the plateau center and edge, respectively. We find that an empirical formula which combines a term motivated by an Efros-Shlovskii (ES) form of variable-range hopping (VRH) [54, 55] with an additional temperature-independent transition rate, $\tau_{s,0}$,
@@
can provide a reasonable fit to the data as shown by the dashed lines in Fig. 4C. Here $\tau_{s,1}^{-1}(T)\sim 1/T$, $T_{ES}$ is the characteristic ES VRH temperature scale, and $\tau_{s,1(0)}$ are prefactors for temperature-(in)dependent transition rates, respectively. While VRH has been used in the past to model longitudinal conduction due to localized states in the quantum Hall effect [55, 56, 57, 58], at the present moment the physical mechanism behind the temperature scaling of $\tau_{s}^{-1}$ here remains nebulous. We comment that there is a wealth of additional information present in the RTN signal, including how the probabilistic weightings of the phase branches (see the right inset of Fig. 1E) evolve with $T$ and $V_{MG}$, that could be analyzed in future studies to inform a microscopic model of the trapping sites present in FQH states.
While we were able to provide evidence that the anyon fluctuations are closely associated with the number of anyon localized states, and that they have consistent temperature scaling properties, it remains unclear whether they arise from device-specific considerations or rather from the fundamental nature of FQH states in monolayer graphene. While fluctuations have not been reported in the same manner in GaAs heterostructure-based fractional interferometry experiments [39, 15, 14] and more recently in bilayer graphene [59], we note that the presence of slower or faster switching timescales could make initial observations and interpretation of anyon fluctuations challenging with standard DC measurement techniques.
@@ Acknowledgments
We thank Abhishek Banerjee, Yuval Ronen, and Steven H. Simon for helpful discussions. The major part of the experiment was supported by DOE (DE-SC0012260). J.R.E. acknowledges support from ARO MURI (N00014-21-1-2537) for sample preparation, measurement, characterization, and analysis. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 20H00354 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. B.I.H. acknowledges support from NSF grant DMR-1231319. Nanofabrication was performed at the Center for Nanoscale Systems at Harvard, supported in part by an NSF NNIN award ECS-00335765.

### v2

@@ Abstract
The search for anyons, quasiparticles with fractional charge and exotic exchange statistics, has inspired decades of condensed matter research. Quantum Hall interferometers enable direct observation of the anyon braiding phase via discrete interference phase jumps when the number of encircled localized quasiparticles changes. Here, we observe this braiding phase in both the $\nu=1/3$ and $4/3$ fractional quantum Hall states by probing three-state random telegraph noise (RTN) in real-time. We find that the observed RTN stems from anyon quasiparticle number $n$ fluctuations and reconstruct three Aharonov-Bohm oscillation signals phase shifted by $32\pi/$ 3, corresponding to the three possible interference branches from braiding around $n$ (mod 3) anyons. Hence, we reveal the anyon braiding phase using new methods that can readily extend to non-abelian states.
@@ Introduction
FQH Fabry-Pérot (FP) interferometers enable direct measurements of the anyon braiding phase. [11, 12]. By partially backscattering current at two quantum point contact (QPC) constrictions, the conductance through the FP cavity modulates with the phase accumulated by edge-traveling quantum Hall (QH) quasiparticles [5, 13]. When the number of localized cavity quasiparticles encircled by the interfering edge changes, the resulting phase shift in the interference signal is equivalent to twice the fundamental exchange phase (modulo $2\pi$) [14, 15]. However, since the observed discrete phase jumps represent charging events in the cavity, resultant Coulomb effects on the interference path can pose complications in probing the braiding effect. In particular, the Coulomb coupling between the bulk and edge must be accounted for when extracting the anyon braiding phase from a small number of phase jump events [16]. Consequently, mitigating Coulomb coupling while maintaining electrostatic tunability has been a major challenge in semiconductor-based interferometers [17, 15].
@@
Figure 1: Graphene FQH interferometer telegraph noise. (A) False-color scanning electron microscope image of a representative device depicting an interference measurement with a single edge channel. Scale bar: $1\mu$m. We measure the conductance $G=I_{d}/(V_{D}^{+}-V_{D}^{-})$ where $I_{d}$ is the measured drain current and $V_{D}^{\pm}$ are the measured voltages at the indicated locations. The top graphite gate is etched into 8 separately tunable regions controlled by voltages $V_{LG}$, $V_{LSG}$, $V_{PG}$, $V_{MG}$, $V_{RSG}$, and $V_{RG}$. Additionally, voltages $V_{LBG/RBG}$ applied to suspended bridges above the two QPCs independently set their transmissions. See Fig. S1 for more details. (B) Hall conductivity $\sigma_{xy}$ and longitudinal conductivity $\sigma_{xx}$ measured with contacts in the left reservoir of the device highlighting the robust plateaus at $\nu=0$, $1/3$, $1$, and $4/3$, which are the quantum Hall states we use in this work. (C) Derivative of $\sigma_{xy}$ showing the expected magnetic field dependence of the well-developed quantum Hall states. (D) Conductance $G$ across the left/right QPC tuned by $V_{LBG/RBG}$ with the right/left QPC deactivated and the bulk in $\nu=4/3$. See Fig. S2 for more details on QPC tunings. Inset: schematics of the edge channel configuration and tunneling at a QPC. As $V_{LBG/RBG}$ increases, the QPC opens more and transmits the outer integer edge (pink) first and eventually transmits the inner fractional edge (blue) with weak backscattering. (E) Three-state RTN in the conductance $G$ when the fractional edge of $\nu=4/3$ is weakly backscattering (corresponding to the tunings marked by the black/purple dots for the left/right QPC in D). Top inset: cartoons of the three possible interference branches from braiding around $n$ (mod 3) anyons, corresponding to the three conductance levels. Right inset: histogram of the conductance over time with 1000 bins. (F) Extracted sinusoidal oscillations from histograms over approximately 8 minutes for each $V_{PG}$ voltage. Each data point shows the central conductance value of a gaussian fit to a histogram peak, shaded by the total counts summed under the gaussian. See Fig. S3 for similar data in $\nu=1/3$. All data is at $T=20$ mK and $B=12$ T unless otherwise noted.
@@
Graphene-based interferometers offer several advantages for realizing anyon braiding, both for the abelian and non-abelian cases. First, both QPCs [18, 19] and interferometers [20, 21, 22, 23, 24] exhibit robust electrostatic tunability for integer QH states. Second, graphite gates reduce Coulomb coupling to reveal Aharonov-Bohm (AB) oscillations [21], which previously required complex quantum well structures in GaAs [17, 15]. Simultaneously, these atomically-flat gates enhance FQH states by screening out charge disorder [25, 26] while allowing precise control of QPC transmission [21, 27, 28] and FP cavity filling factor [29]. Encouragingly, bilayer graphene has displayed a multitude of even-denominator FQH states which may host non-abelian anyons with relatively large energy gaps [30, 31, 32, 33, 34], as required for topologically-protected qubits. Less explored even-denominator states have also been observed in monolayer graphene [35, 36].
@@
Measuring how the magnetic field $B$ changes the total interference phase is crucial in determining if the interferometer falls into the AB or Coulomb-dominated regime [12, 5]. We start for comparison with the $\nu=1$ integer QH state. By repeatedly sweeping the plunger gate, we observe the expected single-sinusoidal oscillations devoid of phase jumps (Fig. 2A). As we change $B$, we observe magnetic-field dependent AB interference (Fig. 2B), where the magnetic field period $\Delta B_{1}$ yields an area $\phi_{0}/\Delta B_{1}=0.80\ \mu\text{m}^{2}$, as expected for interference of electrons in $\nu=1$. A similar measurement in $\nu=1/3$ with varying $B$ and $V_{PG}$ produces an irregular 2D map (Fig. 2C) since repeated plunger gate sweeps are not consistent due to the frequent RTN fluctuations. However, we observe that each plunger gate scan contains stochastic switching between 3 sinusoidal branches, again shifted by $32\pi/$ 3 (Fig. 2D). By superimposing 100 of these sweeps and plotting the histogram of total recorded conductance values over $V_{PG}$, we are able to sample enough instances of the AB oscillation signal of each branch to recreate the three interwoven sinusoids seen in the previous measurement technique discussed in Fig. 1. Fig. 2E exhibits the three interwoven branches at different magnetic fields. As the magnetic field decreases (from the top to bottom panels), we observe that each branch continuously drifts in the increasing $V_{PG}$ direction, creating a negative slope of constant phase (dashed line) analogous to the integer case of Fig. 2B. For $\nu=1/3$, the corresponding flux super-periodicity $\Delta B_{1/3}$ (the magnetic field period for a single branch to return to its initial $V_{PG}$ coordinate) yields $3\phi_{0}/\Delta B_{1/3}=0.83\ \mu\text{m}^{2}$. This value is what one expects for interfering $e/3$ quasiparticles such that $\phi_{0}\rightarrow h/(e/3)=3\phi_{0}$. We similarly observe AB interference and a flux super-periodicity on the fractional edge of $\nu=4/3$ (Fig. S4), while the integer edge shows typical electron AB interference with no observable RTN.
@@
The striking behavior of the anyon branch switching rate and amplitude along the plateau reveals that the quasiparticle fluctuations and coherence are closely correlated with the compressibility of the FQH state. A natural explanation would be to account for the increase in switching rate with an increase in the number of localized states, and thus anyon trapping sites, near the edges of the FQH plateau. However, a simple increase in the individual switching rates of a constant number of traps across the plateau cannot be ruled out with just this observation alone.
@@
In addition to the decrease in visibility, the RTN also shows an increase in switching rate as temperature increases. Fig. 4C shows temperature dependent $\tau_{s}^{-1}$ for two representative data sets, one at the center (red) and the other at the edge (yellow) of the plateau. While switching rate indeed increases quickly at the high temperature limit, there is a dramatic flat saturation of $\tau_{s}^{-1}$ for both data sets at low temperatures, below about $100$ mK and $50$ mK for the plateau center and edge, respectively. We find that an empirical formula which combines a term motivated by an Efros-Shlovskii (ES) form of variable-range hopping (VRH) [55, 56] with an additional temperature-independent transition rate, $\tau_{s,0}^{-1}$,
@@
can provide a reasonable fit to the data as shown by the dashed lines in Fig. 4C. Here $\tau_{s,1}^{-1}(T)\sim 1/T$ and $T_{ES}$ is the characteristic ES VRH temperature scale. While VRH has been used in the past to model longitudinal conduction due to localized states in the quantum Hall effect [56, 57, 58, 59], at the present moment the physical mechanism behind the temperature scaling of $\tau_{s}^{-1}$ here remains nebulous. We comment that there is a wealth of additional information present in the RTN signal, including how the probabilistic weightings of the phase branches (see the right inset of Fig. 1E) evolve with $T$ and $V_{MG}$, that could be analyzed in future studies to inform a microscopic model of the trapping sites present in FQH states.
While we were able to provide evidence that the anyon fluctuations are closely associated with the number of anyon localized states, and that they have consistent temperature scaling properties, it remains unclear whether they arise from device-specific considerations or rather from the fundamental nature of FQH states in monolayer graphene. Based on the uniformity of the AB-modulating sinusoids in Figs. 1F and 2E, we postulate that the anyons are neither hopping from nor getting trapped within regions close to the QPCs, as even small changes in environmental charge would drastically change their tunings and the resultant visibility of the interference signal. Additionally, although fluctuations have not been reported in the same manner in GaAs heterostructure-based fractional interferometry experiments [40, 16, 15] and more recently in bilayer graphene [60], we note that the presence of slower or faster switching timescales could make initial observations and interpretation of anyon fluctuations challenging with standard DC measurement techniques.
@@
Note added: During the preparation of this manuscript, we became aware of a concurrent work using a similar graphene device [64].
@@ Acknowledgments
We thank Abhishek Banerjee, Yuval Ronen, and Steven H. Simon for helpful discussions. The major part of the experiment was supported by DOE (DE-SC0012260). J.R.E. acknowledges support from ARO MURI (N00014-21-1-2537) for sample preparation, measurement, characterization, and analysis. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 20H00354 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. M.E.W. and A.Y. acknowledge support from Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). B.I.H. acknowledges support from NSF grant DMR-1231319. Nanofabrication was performed at the Center for Nanoscale Systems at Harvard, supported in part by an NSF NNIN award ECS-00335765.